OpenStudy (anonymous):

10-2*logx=13 solve for x

jimthompson5910 (jim_thompson5910):

10-2*logx=13 -2*logx=13-10 -2*logx=3 logx=3/(-2) logx = -1.5 x = 10^(-1.5) x = 0.0316227766 So the solution is approximately x = 0.0316227766